The parallel forces F1 = 300 N and F2 = 60 N with the same direction are shown, while A1A2 = l = 100 cm. a) Find the value and point of application for the resultant force. b) Determine the arm d1, when it is known that d2 = 50 cm.


1
Expert's answer
2021-03-09T10:58:46-0500

a)


"300x=(100-x)60\\x=16.7\ cm"

"F=300+60=360\ N"

b)


"300x=(50)60\\x=10\ cm"